OpenStudy (anonymous):

Tamara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 450 customers ordered turkey, 375 customers ordered ham, and 250 customers ordered neither turkey nor ham. Based on these data, how many of the next 1,000 customers will order both turkey and ham? Show your work and use complete sentences.

OpenStudy (tkhunny):

This is not a good question. "How many of the next 1,000 customers will order both turkey and ham?" This question cannot be answered. The following question can be answered... "How many of the next 1,000 customers CAN BE EXPECTED TO order both turkey and ham?" Draw some overlapping circles and sort it out!

OpenStudy (anonymous):

450 + 375 - x + 250 = 1000 1075 - x = 1000 x = 1075 - 1000 x = 75 75 of the next 1000 customers will order both Turkey and Ham. like this?